OpenStudy (moonlitfate):

Simplify (-4+2i)/5i and write in standard form.

OpenStudy (moonlitfate):

I ended up with the answer: \[\frac{ -2+4i}{ 5 }\] It was incorrect, though. I'm not really sure where I went wrong here.

OpenStudy (anonymous):

you just mixed up a sign, I think... how did you start?

OpenStudy (moonlitfate):

I think I started off by separating the whole fraction into two fractions: \[\frac{ -4 }{ 5i } + \frac{ 2i }{ 5i }\]

OpenStudy (moonlitfate):

It's been a bit since I tried to the problem, so I'm trying to remember what to do. I know that the i in the denominator can't be there.

OpenStudy (anonymous):

ah.. What I would have done is to multiply by what my teacher used to call a "well chosen "1"" Multiply by (i / i) which is just one, right?

OpenStudy (moonlitfate):

Hm. I don't really get that. It's been a while since I've taken a traditional math class. But, I think I found what I did wrong from separating the fractions.

OpenStudy (moonlitfate):

\[\frac{ 2i }{ 5i } \] The i cancels out there and it's just 2/5 ; So wouldn't the answer be: \[\frac{ 2 }{ 5 } - \frac{ 4 }{ 5i }\] ?

OpenStudy (anonymous):

\[\frac{ -4 + 2i }{ 5i }\frac{ i }{ i } = \frac{ -4i + 2(i)(i) }{ -5 } = \frac{ -4i-2}{ -5 }\]

OpenStudy (moonlitfate):

Wow, I hate imaginary numbers. :/

OpenStudy (anonymous):

cheer up... they aren't really there ;)

OpenStudy (anonymous):

The trick is often to multiply the top and bottom by something to eliminate the i on the bottom. For a problem with something like (5 + 2i) on the bottom, multiply the top and bottom by (5 - 2i)... then the bottom becomes 25 +10i - 10i -4i^2 = 25 +4 = 29 So you wipe out the i term on the bottom by multiplying by what's called the complex conjugate... just reverse the sign on the i term, in essence.

OpenStudy (anonymous):

btw, the final step to my equations above needs to be to factor out and cancel the negatives on top and bottom... (-4i-2) / (-5) = (-1)(4i+2) / (-1)(5) = (4i + 2) / 5

OpenStudy (moonlitfate):

Thank you. I'll try to remember this, hopefully I'll be able to do similar problems in time for the midterm next week. ;O

OpenStudy (anonymous):

study & practice!!! Good luck!!